Bob's back deck is 12 feet by 10 feet. He wants to make the new area 195 ft.^2 by adding the same dimension to each side. What are the new dimensions?
Answer by Cromlix(4381)

Increase 12feet by x = 12 + x
Increase 10feet by x = 10 + x
(12 + x)(10 + x) = 195
x^2 + 22x + 120 = 195
x^2 + 22x - 75 = 0
(x + 25)(x - 3) = 0
x + 25 = 0
x = -25 (no answer as x <0)
x - 3) = 0
x = 3
New length = 12 + x = 12 + 3 = 15
New width = 10 + x = 10 + 3 = 13
15 feet * 13 feet = 195 ft^2.
Hope this helps.
:-)